OpenStudy (gamereternity):

Which choice lists all the integer solutions of the inequality |x| < 4? A. –4, –3, –2, –1, 0, 1, 2, 3, 4 B. –3, –2, –1, 0, 1, 2, 3 C. –3, –2, –1 D. 0, 1, 2, 3

OpenStudy (gamereternity):

@Chunkymonkay

OpenStudy (chunkymonkay):

hmm.. this one im not sure give me a sec

OpenStudy (chunkymonkay):

I think the answer would be B... but you can tag someone if you want a 100% sure answer.

OpenStudy (gamereternity):

ty

OpenStudy (gamereternity):

@agent0smith

OpenStudy (chunkymonkay):

Np

OpenStudy (gamereternity):

@AloneS

OpenStudy (gamereternity):

@Atsie

OpenStudy (chunkymonkay):

@hartnn

OpenStudy (gamereternity):

@mathmale

OpenStudy (zzr0ck3r):

\(|x|\le 4 \) means \(-4\le x \le 4\) so we want all the integers between \(-4\) and \(4\). Remembers integers are the counting numbers and the negative. i.e. the integers are \(...,-3,-2,-1,0,1,2,3,...\)

OpenStudy (gamereternity):

so b

OpenStudy (zzr0ck3r):

no, is \(-4\le -4\) ?

OpenStudy (gamereternity):

no there equal to

OpenStudy (zzr0ck3r):

so -4 would be in the set. and \(4\le 4\) so \(4\) would also be in it.

OpenStudy (gamereternity):

a right

OpenStudy (zzr0ck3r):

correct